# Mix-LN: Unleashing the Power of Deep Layers by Combining Pre-LN and Post-LN

### v1

@@ 1 Introduction
Recent studies reveal a critical observation regarding the effectiveness of deeper layers in LLMs, particularly those beyond the middle layers. It has been shown that these deeper layers can often be pruned significantly (Yin et al., 2023), or even removed entirely (Gromov et al., 2024; Men et al., 2024), without notably affecting the model’s overall capabilities. Moreover, Li et al. (2024) demonstrated that deeper layers contribute minimally to performance during fine-tuning, further questioning their importance. Unfortunately, this finding has been largely overlooked by the research community, where many see it primarily as an opportunity for model compression (Siddiqui et al., 2024; Zhong et al., 2024; Sreenivas et al., 2024), rather than recognizing it as a potential shortfall in the training process.
@@ 2.1 Layer Normalization and Its Gradient
$\frac{\partial\mathrm{LN}(x^{\prime})}{\partial x^{\prime}}=\frac{1}{\sigma_{x^{\prime}}}I.$ (7)
@@ 4.2 Scaling Up to 7B Model
Evaluating whether the benefits of Mix-LN scale to larger models, such as 7B parameters, is essential to demonstrate if its benefits also hold for larger scale models. To this end, we conducted experiments using the LLaMa-7B architecture, which features an embedding size of 4096 and 32 total layers, following the setup of Zhao et al. (2024). All training configurations were kept identical, with the exception of the layer normalization method. Due to computational constraints, we were able to complete only 13,000 steps of training. Based on our experience, models that exhibit consistent improvements early in training typically retain these advantages through the later stages. The training curve comparison is presented in Figure 4, where it is evident that Mix-LN consistently outperforms Pre-LN during the early training stages of LLaMa-7B.
@@ 4.5 Evaluation with Vision Transformers
To evaluate Mix-LN on non-language models, we replace Pre-LN in ViT models with Mix-LN with $\alpha=0.25$. We train the updated model for 120 epochs on ImageNet-1K following Liu et al. (2022a), using the ConvNeXt (Liu et al., 2022b) configurations. The results clearly demonstrate that the benefits of Mix-LN also generalize to non-language models. The results in Table 4 demonstrate that the benefits of Mix-LN extend to non-language tasks, with performance gains that are more pronounced in larger models (ViT-Small) compared to smaller ones (ViT-Tiny).
@@ 5 Analysis and More Evaluations
6 Related Work
6.1 Normalization in Language Models
To improve the effectiveness of deeper layers in language models, various LN variants have been proposed. For instance, Wang et al. (2019) verified empirically that Post-LN suffers from gradient vanishing in deep Transformers, while Pre-LN facilitates stacking more layers. They consequently introduced dynamic linear combination of layers (DLCL), which connects all previous layers to improve trainability. Similar techniques have been employed in other works (Bapna et al., 2018; Dou et al., 2018). Liu et al. (2020) revealed that Post-LN has strong dependencies on the residual branch, often leading to instability. To address this, Adaptive Model Initialization (Admin) was introduced, which uses additional parameters to control residual dependencies in Post-LN, stabilizing training. DeepNorm (Wang et al., 2024) further improved the trainability of deep Transformers by upscaling the residual connection before applying LN, reducing model updates, and enabling deeper architectures. Additionally, Ding et al. (2021) proposed Sandwich LayerNorm, normalizing both the input and output of each transformer sub-layer. Takase et al. (2022) identified that Post-LN tends to preserve larger gradient norms in deeper layers, potentially leading to more effective training. To address the issue of gradient vanishing in early layers, they introduced B2T, a method that uses a residual connection to bypass all LN except the final one in each layer. We got inspiration from Takase et al. (2022), addressing the limitations of both Pre-LN and Post-LN by combining them. We study Scaled Initialization and Scaled Embed in Appendix B.
6.2 Inefficacy of Deep Layers in LLMs
The Inefficacy of deep layers in LLMs serves as a valid indicator for LLM pruning. Yin et al. (2023) demonstrated that the deeper layers of prominent LLMs like LLaMA and Mistral can be pruned more aggressively than earlier layers, without causing a significant drop in performance. Similarly, Gromov et al. (2024) and Men et al. (2024) further explored layer pruning, identifying the deeper layers of LLMs as typically less essential. Lad et al. (2024) observed that in models like Pythia and GPT-2, deeper layers exhibit strong resilience to interventions, such as layer deletion or swapping. Our work shares similarities with Gromov et al. (2024) in applying angular distance to assess the effectiveness of layers. However, while they identify the inefficacy of deeper layers, they do not offer an explanation for this phenomenon nor propose a solution to address it.
While previous studies often view these characteristics of deeper layers as an opportunity for model compression (Siddiqui et al., 2024; Zhong et al., 2024; Sreenivas et al., 2024), we argue that this behavior reveals a deeper training shortfall—primarily due to the widespread use of Pre-LN. In response, we introduce Mix-LN, a novel method that enhances the effectiveness of deeper layers, ensuring that the entire architecture is more effectively trained and fully leverages the network’s parameters.

### v2

@@ 1 Introduction
Recent studies reveal a critical observation regarding the effectiveness of deeper layers in LLMs, particularly those beyond the middle layers (Jaiswal et al., 2024; Sun et al., 2025; Liu et al., 2022b). It has been shown that these deeper layers can often be pruned significantly (Yin et al., 2023), or even removed entirely (Gromov et al., 2024; Men et al., 2024), without notably affecting the model’s overall capabilities. Moreover, Li et al. (2024) demonstrated that deeper layers contribute minimally to performance during fine-tuning, further questioning their importance. Unfortunately, this finding has been largely overlooked by the research community, where many see it primarily as an opportunity for model compression (Siddiqui et al., 2024; Zhong et al., 2024; Sreenivas et al., 2024), rather than recognizing it as a potential shortfall in the training process.
@@ 2.1 Layer Normalization and Its Gradient
$\frac{\partial\mathrm{LN}(x^{\prime})}{\partial x^{\prime}}\approx\frac{1}{\sigma_{x^{\prime}}}I.$ (7)
@@ 4.2 Scaling Up to 7B Model
It is essential to evaluate if the benefits of Mix-LN also hold for larger model sizes. To this end, we conducted experiments using the LLaMa-7B architecture, which features an embedding size of 4096 and 32 total layers. All training configurations were kept identical, except for the layer normalization method. Due to computational constraints, we managed to complete only 13,000 steps of training. The training curve is presented in Figure 4. We can see that it is possible to scale the gain of Mix-LN to larger-scale training.
Nevertheless, we found that Mix-LN becomes more sensitive to $\alpha$ when scaling up to 7B training. In general, smaller values of $\alpha$ and a longer warm-up period are required to stabilize the training of Mix-LN at a larger scale. For instance, we use a relatively small $\alpha=6.25\%$ here to ensure stable training of LLaMa-7B. We conjecture that this instability is relevant to Post-LN, as using Post-LN alone also leads to the same issue. Addressing this challenge presents an interesting research direction.
Table 2: Pre-training Perplexity of OLMO-1B and OLMO-7B.
Model #Params Pre-LN Pre-LN + LNS
LLaMA-1B 1B 23.14 21.99
LLaMA-7B 7B 16.06 14.98
Table 3: Pre-training Perplexity of Qwen2.5-0.5B.
Model #Params Pre-LN Pre-LN + LNS
Qwen2.5-0.5B 0.5B 20.62 19.57
@@ 4.5 Evaluation with Vision Transformers
To evaluate Mix-LN on non-language models, we replace Pre-LN in ViT models with Mix-LN with $\alpha=0.25$. We train the updated model for 120 epochs on ImageNet-1K following Liu et al. (2022a), using the ConvNeXt (Liu et al., 2022c) configurations. The results clearly demonstrate that the benefits of Mix-LN also generalize to non-language models. The results in Table 6 demonstrate that the benefits of Mix-LN extend to non-language tasks, with performance gains that are more pronounced in larger models (ViT-Small) compared to smaller ones (ViT-Tiny).
@@ 5 Analysis and More Evaluations
6 Limitations
One limitation of Mix-LN is the introduction of Post-LN leads to instability issues when scaling up training, i.e., $>$7B. We conjecture that this instability is relevant to Post-LN, as using Post-LN alone also leads to the same issue. An interesting research direction is to explore how to mitigate the ineffectiveness of later layers without relying on Post-LN or to address the instability associated with Post-LN.
@@ 7 Conclusion
8 Acknowledgment
This publication is partly financed by the Dutch Research Council (NWO) under the grant no. NWO-2023.027/L1.
@@ Appendix B Compatibility to Advanced
Appendix C Related Work
C.1 Normalization in Language Models
To improve the effectiveness of deeper layers in language models, various LN variants have been proposed. For instance, Wang et al. (2019) verified empirically that Post-LN suffers from gradient vanishing in deep Transformers, while Pre-LN facilitates stacking more layers. They consequently introduced dynamic linear combination of layers (DLCL), which connects all previous layers to improve trainability. Similar techniques have been employed in other works (Bapna et al., 2018; Dou et al., 2018). Liu et al. (2020) revealed that Post-LN has strong dependencies on the residual branch, often leading to instability. To address this, Adaptive Model Initialization (Admin) was introduced, which uses additional parameters to control residual dependencies in Post-LN, stabilizing training. DeepNorm (Wang et al., 2024) further improved the trainability of deep Transformers by upscaling the residual connection before applying LN, reducing model updates, and enabling deeper architectures. Additionally, Ding et al. (2021) proposed Sandwich LayerNorm, normalizing both the input and output of each transformer sub-layer. Takase et al. (2022) identified that Post-LN tends to preserve larger gradient norms in deeper layers, potentially leading to more effective training. To address the issue of gradient vanishing in early layers, they introduced B2T, a method that uses a residual connection to bypass all LN except the final one in each layer. Huang et al. (2025) unveiled that LN layers suffer from loss spikes and gradient spikes during LLM training. We got inspiration from Takase et al. (2022), addressing the limitations of both Pre-LN and Post-LN by combining them. We study Scaled Initialization and Scaled Embed in Appendix B.
C.2 Inefficacy of Deep Layers in LLMs
The Inefficacy of deep layers in LLMs serves as a valid indicator for LLM pruning. Yin et al. (2023) demonstrated that the deeper layers of prominent LLMs like LLaMA and Mistral can be pruned more aggressively than earlier layers, without causing a significant drop in performance. Similarly, Gromov et al. (2024) and Men et al. (2024) further explored layer pruning, identifying the deeper layers of LLMs as typically less essential. Lad et al. (2024) observed that in models like Pythia and GPT-2, deeper layers exhibit strong resilience to interventions, such as layer deletion or swapping. Our work shares similarities with Gromov et al. (2024) in applying angular distance to assess the effectiveness of layers. However, while they identify the inefficacy of deeper layers, they do not offer an explanation for this phenomenon nor propose a solution to address it. Building on the insights from this paper, Sun et al. (2025) introduced the concept of the “Curse of Depth” to highlight the pitfalls of deeper layers in LLMs. To address this issue while avoiding the instability associated with Post-LN, they proposed LayerNorm Scaling to scale down the output of LN by its depth.